the substation bus bar is made up of 2-inches round copper bars, 20- feet long. what is the resistance of each bar if the resistivity is 1.7241 x 10 -6 ω-cm.
1
Expert's answer
2021-09-26T21:01:39-0400

Length (l)="\frac{20}{3.281}=6.096m"

r="2\times0.0254=0.0508m"

"R=\frac{\rho l}{A}"

Put value

"A=\pi(0.0508)^2=8.11\times10^{-3}m^2"

"\rho=1.7241\times10^{-8}\Omega\times m"

"R=\frac{\rho l}{A}"



"R=\frac{1.7241\times10^{-8}\times6.096}{3.14\times0.0508^2}\\R=\frac{1.7241\times10^{-8}\times6.096}{8.11\times10^{-3}}=5.185\times10^{-5}\Omega"